# Categorical Reparameterization with Denoising Diffusion models

### v1

@@ 4 Experiments
In this section, we conduct an empirical evaluation of our method on several benchmark problems, including polynomial programming, Sudoku problem solving, and applications to variational inference and generative modeling. We compare our method against three representative baselines from the literature: the Straight-Through (ST) estimator (bengio2012estimating), the Gumbel-Softmax estimator (more precisely, its Straight-Through variant) (jang2017categorical), and the more recent ReinMax method (liu2023reinmax). Among these, ReinMax reports state-of-the-art performance on most of the benchmarks it considers and, to the best of our knowledge, is one of the most recent approaches addressing the same class of problems as ours. For this reason, and since liu2023reinmax show that ReinMax consistently outperforms several earlier alternatives, we do not include additional baselines in our comparison. All the hyperparameters are detailed in Table 5. For ReDGE and its variants we use the linear schedule $(\alpha_{t},\sigma_{t})=(t,1-t)$ and ${t_{k}}=k/(n-1)$ for $k\in[0:n-1]$ where $n$ is the number of diffusion steps. We use ReDGE-Max with $\pi_{1}=\mathcal{N}(0,\mathbf{I}_{K})^{\otimes L}$ as initialization. Finally, we use the hard gradient estimator for each method.
4.1 Polynomial programming
We illustrate our approach on the polynomial programming toy problem also considered in tucker2017rebar; grathwohl2018relax; paulus2020raoblackwell; liu2023reinmax. The factorized distribution $\pi_{\!\theta}$ is such that for all $i\in[L]$, $\pi^{i}_{\!\theta}=\text{Bernoulli}(\frac{\exp(\theta^{i2})}{\exp(\theta^{i1})+\exp(\theta^{i2})})$ with $\theta\in{\mathbb{R}}^{L\times 2}$. Following prior works, we consider a fixed target vector $c=0.45$ and solve the optimization problem
$\min_{\theta\in{\mathbb{R}}^{L\times 2}}\frac{1}{L}\mathbb{E}_{\pi_{\!\theta}}\left[\|X\!\cdot\!(0\,\,\,1)^{\top}-c\!\cdot\!\mathbf{1}_{L}\|_{p}^{p}\right],$ (16)
for a fixed exponent $p\geq 1$ and with $L=128$. The optimal policy is the one that puts all the mass on the matrix with $e_{1}$ in each row. We report results on this benchmark in Fig. 2. We also emphasize several limitations of this example, which to our knowledge have not been explicitly discussed in the gradient-estimation literature and somewhat undermine its relevance as a stand-alone evaluation:
(1).  The objective is separable and identical across dimensions, so the gradient can be recovered from only two loss evaluations (one per coordinate value).
(2).  ST estimator performs poorly in this experiment. However, note that discrete objective is determined entirely by the values of $f$ at the vertices of the product simplex. Consequently, any extension on ${\mathbb{R}}^{L\times 2}$ that matches $f$ on these vertices defines the same discrete problem, yet may induce a very different optimization landscape. As an illustration, consider the extension
$f:\ x\in\mathbb{R}^{L\times 2}\ \mapsto\ \frac{1}{L}\sum_{i=1}^{L}(-c)^{p}x^{i1}+(1-c)^{p}x^{i2}\;,$
which is linear and coincides with (16) on the vertices. For this relaxation, hard ST yields a low-variance unbiased gradient estimator (and soft ST yields the exact gradient) that performs well. See Appendix D.3 for results with this linear relaxation.
(3).  Finally, ReinMax is based on a second-order Taylor approximation of $f$. Hence, when $p=2$ in (16) the estimator is exact (proof in Appendix B.2); for other degrees it is no longer exact, though it often remains a close approximation in practice. This exactness is specific to quadratic objectives and does not extend to general $f$.
4.2 Variational Inference for Gaussian Mixture Model
As a second benchmark, we follow the Gaussian mixture variational inference experiment of liu2019rao, framed in the setting of blei2017variational.
Generative model.
We consider a $d$-dimensional Gaussian mixture model with $K$ components. In our experiments we take $d=2$, $K=20$, and draw $N=500$ observations. For $\pi\in\Delta^{K-1}$, the generative model consists in drawing i.i.d. cluster assignments $Z\coloneqq(Z^{1},\dotsc,Z^{N})\overset{\mathrm{i.i.d.}}{\sim}p_{z}\coloneqq\mathrm{Categorical}(\pi)$, then we draw a matrix $M\in{\mathbb{R}}^{K\times d}$ i.i.d. wth i.i.d. rows $M^{i}\sim p_{m}\coloneqq\mathcal{N}(0,\sigma_{0}^{2}\,\mathbf{I}_{d})$ for $i\in[K]$. Finally, we draw a matrix of $N$ observations $Y\in{\mathbb{R}}^{N\times d}$ with $Y^{i}\sim p_{y}(\cdot|M,Z^{i})\coloneqq\mathcal{N}(M^{Z^{i}},\sigma_{y}^{2}\,\mathbf{I}_{d})$.
Variational family.
Exact Bayesian inference over the means and cluster assignments $(M,Z)$ given a realization $Y=y$ is intractable, so we approximate the posterior with a mean-field variational family. Following the setup in liu2019rao, we consider the variational family
$q_{\phi}(\mathrm{d}z,\mathrm{d}m)=\pi_{\!\theta}(\mathrm{d}z)\prod_{i=1}^{K}\delta_{\hat{m}^{k}}(\mathrm{d}m^{k})\;.$
where $\phi\coloneqq(\theta,\hat{m}^{1},\dotsc,\hat{m}^{K})$ and $\pi_{\!\theta}$ is the factorized categorical distribution (2) over the cluster assignments with $L=N$ and we have implicitely replaced the cluster assignments with their one-hot encodings. Since the mean component is degenerate, rather than minimizing the KL which would require absolute continuity, we instead minimize the following objective which can be seen as a variational objective over the cluster assignment combined with a MAP estimation over the mean components;
$F(\theta;\hat{m}^{1:K})\coloneqq\sum_{i=1}^{N}\mathbb{E}_{\pi^{i}_{\!\theta}}\big[\log\pi^{i}_{\!\theta}(Z^{i})-\log p_{y}(y^{i}|\hat{m},Z^{i})\\ \textstyle-\log p_{z}(Z^{i})\big]-\sum_{k=1}^{K}\log p_{m}(\hat{m}^{k})\;.$
The results are reported in Table 1. Overall, ReDGE-Cov converges substantially faster and reaches the best final ELBO among all methods. In contrast, ReDGE-Max tracks the performance of ReinMax, while Gumbel-Softmax remains behind the vanilla ReDGE baseline both in terms of final value and speed.
We show in table 1 the final ELBO of each sampler, averaged over the 100 last iterations (with the standard deviation) as well as the final clustering accuracies.
Table 1: Gaussian Mixture Model variational inference results. We report the mean and standard deviation of the negative ELBO (NELBO) over the final 100 training iterations (lower is better), as well as the final clustering accuracy (higher is better).
Sampler Final NELBO (mean $\pm$ std) Clustering accuracy (mean $\pm$ std)
Gumbel-Softmax $1296.11\pm 87.66$ $0.56\pm 0.07$
Straight-Through $4380.48\pm 126.67$ $0.45\pm 0.04$
ReinMax $1175.73\pm 78.78$ $0.62\pm 0.10$
ReDGE $1716.58\pm 127.46$ $\mathbf{0.64\pm 0.10}$
ReDGE-Max $1186.59\pm 58.98$ $0.61\pm 0.09$
ReDGE-Cov $\mathbf{1040.04\pm 97.89}$ $0.60\pm 0.09$
4.3 Sudoku
We consider partially solved sudoku grids and frame their completion as an optimization problem. We parameterize a factorized categorical law $\pi_{\!\theta}$ over the sudoku grid with each cell represented by a categorical distribution; i.e. $\pi_{\!\theta}$ is a distribution over $\mathsf{V}^{81}$ with $V$ the set of one-hot encodings of length $9$. Each row/column/block denoted $g$ is represented by a set of indexes $i\in[81]$, and we define the digit count function $s_{g}:X\in\mathsf{V}^{81}\mapsto\sum_{i=1}^{81}\mathds{1}_{i\in g}X^{i}$, which outputs the all-ones vector exactly when the digits in $g$ form a valid permutation (i.e., each digit appears once). We use a quadratic penalty as a relaxed violation count and optimize its expectation under $\pi_{\!\theta}$: $F(\theta)\coloneqq\mathbb{E}_{\pi_{\!\theta}}[\sum_{g}\big\|s_{g}(X)-\mathbf{1}_{9}\big\|_{2}^{2}]\;.$ At first sight, taking $\pi_{\!\theta}$ to be fully factorized across cells may seem too restrictive, since valid Sudoku grids exhibit strong dependencies. The key point, however, is that while $\pi_{\!\theta}$ is mean-field conditional on a fixed $\theta$, the learning dynamics are not: the loss is highly non-separable, and each stochastic gradient step updates many cell logits jointly through shared row/column/block constraints. Consequently, dependencies are introduced through the optimization procedure itself. During training, updates are computed from random samples of the grid. Therefore the parameter iterate is itself random: after $T$ steps, $\theta_{T}$ is a random variable defined by the stochastic recursion induced by the optimizer. The distribution of an output grid produced after $T$ steps from initialization $\theta_{0}$ is thus the mixture $\mathbb{E}\!\left[\pi_{\!\theta_{T}}|\theta_{0}\right]$. Although each component in this mixture factorizes, the mixture does not: the shared optimization noise couples all coordinates through the non-separable constraints, allowing the resulting predictor to place most of its mass on globally consistent Sudoku configurations despite the mean-field parameterization.
Table 2 reports the expected quadratic constraint-violation objective (lower is better) together with the solved rate (zero violations). The diffusion-based estimators consistently improve over standard baselines: ReDGE-Cov attains the lowest mean loss, and ReDGE achieves the highest solved rate, outperforming the remaining baselines. Notably, Straight-Through yields substantially higher losses and variance, suggesting unstable optimization on this highly non-separable penalty. Among our variants, the covariance correction provides the most reliable decrease in the penalty (best mean), while ReDGE slightly trades off penalty minimization for a higher probability of reaching an exactly feasible grid (best solved %).
Table 2: Sudoku results. We report the mean and standard deviation of the loss across runs, as well as the percentage of solved Sudokus (zero violations).
Sampler Loss (mean $\pm$ std) Solved (%)
Gumbel-Softmax 14.46 $\pm$ 10.67 15.42
Straight-Through 31.29 $\pm$ 23.47 9.07
ReinMax 13.21 $\pm$ 9.47 18.31
ReDGE 9.94 $\pm$ 8.63 22.22
ReDGE-Max 11.27 $\pm$ 8.57 14.1
ReDGE-Cov 8.18 $\pm$ 7.13 20.76
4.4 Categorical VAE
Table 3: Categorical VAE results on MNIST for different latent and categorical dimensionalities. We report the final average loss (mean $\pm$ standard deviation) across runs.
Sampler L=32,K=64L=32,\,K=64 L=30,K=10L=30,\,K=10 L=48,K=2L=48,\,K=2
Gumbel-Softmax $85.37\pm 0.79$ $79.33\pm 0.50$ $88.23\pm 0.21$
Straight-Through $106.17\pm 0.03$ $101.51\pm 0.05$ $99.22\pm 0.10$
ReinMax $86.42\pm 1.50$ $80.98\pm 0.35$ $87.61\pm 0.25$
ReDGE $101.90\pm 0.49$ $89.48\pm 0.33$ $87.40\pm 0.21$
ReDGE-Max $86.46\pm 0.67$ $80.20\pm 0.68$ $87.76\pm 0.12$
ReDGE-Cov $\mathbf{81.99\pm 0.06}$ $\mathbf{78.80\pm 0.04}$ $\mathbf{82.01\pm 0.08}$
We train a Bernoulli VAE on binarized MNIST (kingma:welling:2014; rezende2015variational) following the setups in tucker2017rebar; grathwohl2018relax; liu2023reinmax. The encoder is a neural network that maps the input image $x\in\{0,1\}^{784}$ to logits $\varphi_{\theta}(x)\in{\mathbb{R}}^{L\times K}$ and defines the mean-field posterior $\pi_{\!\theta}(\cdot|x)$ (2). The decoder is a neural network modeling pixel logits $\eta_{\phi}(z)\in{\mathbb{R}}^{784}$ given a latent $z\in{\mathbb{R}}^{L\times K}$ to produce the decoding distribution $p_{\phi}(\cdot|z)=\prod_{j=1}^{784}\text{Bernoulli}\big(\sigma(\eta_{\phi}(z)^{j})\big)$ where $\sigma$ is the sigmoid function. Given a dataset $(X_{i})^{N}_{i=1}$, we optimize jointly in $(\theta,\phi)$ the negative ELBO,
$F(\theta;\phi)\coloneqq\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}_{\pi_{\!\theta}(\cdot|X_{i})}\big[\log p_{\phi}(X_{i}|Z)\big]\\ -\textstyle\text{KL}\big(\pi_{\!\theta}(\cdot|X_{i})\,\big\|\,p_{z}\big)\;,$
where $p_{z}\coloneqq\text{Uniform}(\mathsf{X})$ is the discrete uniform distribution over $\mathsf{X}$. The results are summarized in Table 3 for three different configurations of $(L,K)$. ReDGE-Cov achieves a better final loss as well as a faster convergence in all the settings we considered (see more results in Section D.3), outperforming all the other baselines.
Figure 3: Categorical VAE training curves for the configuration with L=48,K=2L=48,\,K=2.
Runtime.
Our method introduces extra computation from the number of diffusion steps used. However, like Straight-Through and Gumbel-Softmax, each gradient estimate requires only a single evaluation of the objective $f$, unlike REBAR/RELAX and related variance-reduction methods which typically require multiple evaluations (tucker2017rebar; grathwohl2018relax; liu2019rao). The cost incurred by the number of diffusion steps is an $f$-independent additive overhead and is negligible in our settings where evaluating $f$ (e.g., a network forward pass or constraint computation) dominates runtime. Hence, our runtime is comparable to the competitors, since we use a small number of diffusion steps (typically $3$ to $5$). Runtime measurements are provided in Table 4.
Table 4: Average runtime per training epoch for the Categorical VAE experiment with L=48,K=2L=48,\,K=2. We report the mean and standard deviation over epochs (in seconds).
Sampler Time per epoch (s, mean $\pm$ std)
Gumbel-Softmax $5.16\pm 0.07$
Straight-Through $5.20\pm 0.20$
ReinMax $5.39\pm 0.15$
ReDGE $5.80\pm 0.32$
ReDGE-Max $6.64\pm 0.21$
ReDGE-Cov $6.15\pm 0.24$

### v2

@@ 4 Experiments
In this section, we evaluate our method on benchmark problems spanning Sudoku solving and generative modeling. Further experiments are given in Appendix E and F. We compare against three representative baselines: the Straight-Through (ST) estimator (Bengio et al., 2013), Gumbel-Softmax (using its straight-through variant) (Jang et al., 2017), and ReinMax (Liu et al., 2023a). We focus on these since ReinMax is a recent strong method that reports state-of-the-art results and is shown to outperform several earlier alternatives (Liu et al., 2023a), so we omit additional baselines. In addition, we don’t compare against REBAR/RELAX-style estimators (Tucker et al., 2017; Grathwohl et al., 2018) because they are meta-estimators that wrap a base estimator with learned control variates and additional tuning. Our method could in principle be used as the base reparameterization within these frameworks, which we leave to future work.
All hyperparameters are reported in Section G.1. We also report the runtime and memory usage in Section G.4. For all methods we use the hard version. For ReDGE and its variants, we use the linear schedule $(\alpha_{t},\sigma_{t})=(1-t,t)$ (Lipman et al., 2023; Esser et al., 2024). For the timesteps we first specify $t_{1}$} and then set ${t_{k}}={t_{1}}+(1-{t_{1}})k/(n-1)$ for $k\in[2:n-1]$.
4.1 Inference-time guidance with Masked Diffusion
We start by providing some necessary background on Masked Diffusion models (MDM) (Shi et al., 2024; Sahoo et al., 2024). We defer a more formal introduction to Appendix D.
Masked diffusion.
Let $p$ be a target distribution defined on $\mathsf{X}$ with the vocabulary augmented by the mask token m. MDMs provide an approximate sampler for $p$ via an iterative unmasking process. We denote the learned model distribution by $p^{\mathsf{d},\psi}_{0}$, where $\psi$ are the model parameters. We use the superscript $\mathsf{d}$ (for discrete) to avoid conflict with the Gaussian diffusion notation. MDMs rely on a clean-data predictor $p^{\mathsf{d},\psi}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}k}(\cdot|X_{k})$ that outputs a factorized categorical approximation (2) to the posterior of $X_{0}\sim p$ given a partially masked state $X_{k}$, under a joint distribution where $X_{k}$ is obtained from $X_{0}$ by setting independently across the dimensions $X^{i}_{k}=\texttt{m}$ with probability $(1-\beta_{k})$ and $X^{i}_{k}=X^{i}_{0}$ otherwise. $(\beta_{k})$ is chosen as a decreasing schedule. Sampling proceeds by simulating a Markov chain $(X_{0:M})$ where $X^{i}_{M}=\texttt{m}$ for all $i\in[K]$, and given $X_{k}$, we first draw an approximate solution $\hat{X}_{0}\sim p^{\mathsf{d},\psi}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}k}(\cdot|X_{k})$, then sample $X_{k-1}$ by keeping the unmasked entries of $X_{k}$ fixed and, for masked entries, setting $X_{k-1}^{i}=\hat{X}_{0}^{i}$ with probability $(\beta_{k-1}-\beta_{k})/(1-\beta_{k})$ (otherwise $X_{k-1}^{i}=\texttt{m}$).
Inference-time guidance.
Given a reward $r$ we want to steer sampling at test time by locally modifying the model’s step-wise predictive distribution to favor samples with higher reward. Following (Murata et al., 2025), this can be achieved by training, given $x_{k}$ at diffusion step $k$, a factorized variational distribution $\pi_{\!\theta}(\cdot|x_{k})$ to approximate the tilted distribution with p.m.f. at $x_{0}$ proportional to $\smash{\exp(-r(x_{0}))p^{\mathsf{d},\psi}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}k}(x_{0}|x_{k})}$. This is done by minimizing the forward KL divergence objective
$F_{k}(\theta)\!\coloneqq\!\mathbb{E}_{\pi_{\!\theta}(\cdot|x_{k})}[r(X_{0})]+\mathrm{KL}(\pi_{\!\theta}(\cdot|x_{k})\|p^{\mathsf{d},\psi}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}k}(\cdot|x_{k}))\;.$ (13)
We then draw $\hat{X}_{0}$ from the obtained proposal and then sample $X_{k-1}$ as previously done. We provide more details in Appendix D. We consider two such applications in the next subsections. In all cases, we optimize the logits directly by setting $\varphi_{\theta}=\theta$ and treating $\theta\in\mathbb{R}^{L\times K}$ as the optimization variable. We detail the guidance algorithm in Appendix D.
4.1.1 MDM Guidance for solving Sudoku puzzles.
We follow Ye et al. (2024) and train a masked diffusion model (MDM) to approximate the distribution $p(\cdot|\mathbf{c})$ over valid completions of an incomplete Sudoku grid $\mathbf{c}$, viewed as a categorical distribution on $\mathsf{V}^{81}$, where $\mathsf{V}$ denotes the set of one-hot vectors of length $10$ and the mask m is $e_{10}$. Let $\mathcal{G}$ denote the 27 constraint groups (rows, columns, and blocks). For $g\in\mathcal{G}$, define the digit-count map $s_{g}(X)\coloneqq\sum_{i\in g}PX^{i}$, where $P$ drops the mask coordinate so that $s_{g}(X)\in\mathbb{R}^{9}$ counts digits in $g$. We use the reward $r(x)\coloneqq\sum_{g\in\mathcal{G}}\bigl\|s_{g}(x)-\mathbf{1}_{9}\bigr\|_{2}^{2}.$
Figure 2: Masked diffusion guidance on Sudoku: fraction solved and mean constraint violations (1000 test puzzles, 10 seeds, 20 diffusion steps), for early (1%) and late (90%) checkpoints, as a function of the gradient-step budget. For each estimator, we sweep hyperparameters and learning rates, select the setting that minimizes the AUC of mean violations over step budgets (100–2000), and plot its violations and solve rate across budgets.
For the first experiment, we use two checkpoints with very different baseline performance: an early model that solves about $1\%$ of the $1000$ test Sudokus and a late model that solves $90\%$. We apply inference-time guidance by optimizing (13), estimating the reward-gradient term with hard gradient estimators and differentiating the KL term exactly.
Results.   The results are given in Figure 2. From the 1% checkpoint, guidance with ReDGE raises the solve rate to 89%, outperforming the strongest baselines, which plateau in the mid-80s. Straight-Through performs substantially worse, and using a smaller $t_{1}$} (ReinDGE) improves over the larger-$t_{1}$} ReinMax special case. Starting from the 90% checkpoint, guidance further improves performance, reaching solve rate 93% with ReDGE-Cov.
Figure 3: Solving Sudoku without pre-trained MDM. Similarly to Figure 2 we select a single configuration by minimizing the area under the mean-violation curve over budgets (100 to 8e4 steps)
Direct optimization without pre-training.
We also study a no-prior variant in which we drop the MDM entirely and directly optimize $\mathbb{E}_{\pi_{\!\theta}}[r(X)]$ w.r.t. the parameters of factorized categorical distribution $\pi_{\!\theta}$. The Sudoku clues in $\mathbf{c}$ are enforced by setting the logit of the observed digit to a very large value at each clue location after every gradient step (whereas with a pre-trained MDM this conditioning is already reflected in the posterior initialization). Surprisingly, the best-performing estimators achieve solve rates in the mid-to-high 90s, substantially higher than what is obtained when guiding from the weak 1% checkpoint, while ReDGE and Gumbel-Softmax lag behind.
Figure 4: Average violations heatmap as a function of $n$ and the timestep $t_{1}$}. For each configuration $(n,{t_{1}})$ we report the lowest AUC obtained over a sweep of four learning rates.
Finally, we provide a comprehensive heatmap summarizing how the performance of ReDGE and ReDGE-Cov behaves as a function of $(n,{t_{1}})$. We make three empirical remarks; (i) we can see that in all cases a smaller $t_{1}$} results in worst performance, as suggested by our theoretical analysis and Figure 1. (ii) Despite its strong performance ReDGE-Cov is more sensitive to the hyperparameters than ReDGE. (iii) Using more diffusion steps doesn’t affect the performance much, except for very small $t_{1}$. This connection between $t_{1}$, $n$ and the performance of our gradient estimators is discussed and studied in more detail in Appendix E.2.
4.1.2 Reward-guided image generation
We next apply inference-time guidance to discrete image generation with a class-conditional pretrained MaskGIT (Chang et al., 2022) model trained on the ImageNet dataset and operating on VQ-VAE codes. We generate images at resolution $384\times 384\times 3$ (Besnier et al., 2025) by sampling a sequence of discrete latent codes $K]^{L}$, where each image is represented by $L=576$ codes, each taking one of $K=16384$ codebook entries. Each image is thus represented by a latent embedding in ${\mathbb{R}}^{576\times d}$ with $d=8$. We write $E$ for the embedding matrix in ${\mathbb{R}}^{K\times d}$. Given $x\in\mathsf{X}$, the reward consists in decoding the embedding $x\cdot E\in{\mathbb{R}}^{L\times d}$ and then computing the CLIP score (Radford et al., 2021; Hessel et al., 2021) with a target prompt.
Figure 5: Left: Average CLIP score for CLIP-guided image generation. Middle and right: sensitivity of CLIP score to the temperature parameters $t_{1}$} and $\tau$. We report the mean over 200 images; for each estimator, we sweep hyperparameters and learning rates, select the setting that maximizes the AUC of CLIP score over gradient steps budget.
Figure 6: ReDGE samples generated by CLIP-guided MaskGIT from the prompts shown.
Results.   As shown in Figure 5, guidance monotonically improves CLIP score as a function of the gradient step budget. ReDGE, ReDGE-Cov, and Straight-Through achieve comparable performance, outperforming ReinDGE and Gumbel-Softmax and substantially surpassing ReinMax. The middle panel of Figure 5 shows that performance peaks at ${t_{1}}=0.9$, a regime that is closer to straight-through behavior and is consistent with Straight-Through also performing well. However, strong performance persists for ${t_{1}}\in\{0.5,0.7\}$, indicating a broad operating range for ReDGE and suggesting that its gains are not solely driven by the near-Straight-Through limit. On the other hand, ReDGE-Cov and Gumbel-Softmax exhibit substantially higher hyperparameter sensitivity.
Takeaways and practical tuning.
Across our benchmarks, ReDGE is the most reliable, delivering strong performance across diverse objectives, whereas our variants as well as Straight-Through, ReinMax, and Gumbel-Softmax are more setting-dependent. ReDGE-Cov can be particularly strong in favorable regimes but is more hyperparameter-sensitive than ReDGE, which offers the best robustness–performance trade-off. For tuning, a small number of diffusion steps (e.g., $n=3,5$) coupled with a moderate $t_{1}$ (e.g. $t_{1}\in\{0.5,0.7,0.9\}$) is a strong default. The endpoint $t_{1}=1$ recovers the straight-through limit and is a useful reference when Straight-Through is competitive, yet strong results often persist for intermediate $t_{1}$, indicating gains beyond the near-Straight-Through regime.